# Causal3D: A Comprehensive Benchmark for Causal Learning from Visual Data

### v1

@@ 4.4 Causal Discovery from Few Images
Evaluated Methods. Unlike traditional causal discovery methods that rely on tabular data or large-scale training images, we leverage pretrained VLMs to perform causal discovery using a small number of images. In this section we use ChatGPT [^2] https://platform.openai.com/docs/api-reference/introduction, Gemini [^3] https://ai.google.dev/api?lang=python and Claude [^4] https://www.anthropic.com/api to discover causal relations by few images examples and textual prompts in real and hypothetical scenes. The VLMs are tasked with uncovering causal relations by generating adjacency matrices representing causal graphs. Each prompt specifies key variables in the scenes, explicitly guiding the models to infer causal structures. We use 4 different prompting strategies, as detailed in Tab. 3 in the Appendix B.
1) Basic Prompts: General instructions designed to broadly guide the models in identifying causal relations.
2) Explicit Function Prompts: The model is designated as a causal discovery expert and directed to identify precise causal relations among image variables.
@@
4) Few-Shot: The model is provided with 3 exemplar cases of causal discovery before performing the task. This strategy helps the model generalize better by leveraging prior examples to improve its identification of causal relations. Performance was evaluated by comparing the generated adjacency matrices with the ground truth, providing a quantitative measure of the causal discovery task.
Figure 5: Realistic scenes.
Figure 6: Hypothetical scenes.
Figure 7: Causal discovery results of VLMs in various scenes, averaging over datasets with 2, 3, 4, 5 nodes in causal graph.
Table 2: Causal intervention in VLMs: We use 3 real-world rules to evaluate the intervention tasks (Acc. (%) is reported).
Models Reflection Convex Lens Magnetic Field
@@
Results and Analysis. Demonstrated in Fig. 7, models performed significantly better in real-world settings (e.g., governed by Hooke’s Law, magnetic fields and etc.) compared to hypothetical scenes, benefiting from prior knowledge of physical laws embedded in their LLM components. This prior knowledge guarantees strong performance in causal discovery tasks. However, as causal relations became more complex and involved more variables, performance declined, highlighting the challenge of uncovering causality in such scenes. In the hypothetical scenes, the results, as shown in Fig. 7, reveal poor performance. This suggests that when prior real-world knowledge is unavailable and only a limited number of images are provided, models consistently fail to uncover latent causal relations. Consequently, current closed-source VLMs have been shown to be unreliable for causal discovery in hypothetical settings.
Views and Backgrounds. Causal3D offers multi-view images of the same scene with both virtual and realistic backgrounds, enabling in-depth analysis of how different viewpoints and background contexts impact causal discovery performance. We show the case study in Spring, and Parabola scenes in Fig. 8. Within the experiments, different 3D views affect the performance in causal discovery. Interestingly, our results reveal that intuitive views, such as a front view, are not always the most effective for uncovering latent causality. Different viewpoints can either increase or decrease task difficulty, with no single “golden standard” view for causal discovery. Additionally, when comparing multi-view and single-view inputs in the causal discovery task, multi-view performance varies depending on the scene (as shown in Fig. 11). In scenes with simple causal relations (e.g., a spring system with a linear relation among three variables), multi-view inputs tend to degrade model performance, possibly introducing unnecessary noise. Conversely, in more complex 3D scenes with intricate causal dependencies, like a parabola scene involving nonlinear relations among four variables in the virtual background, multi-view perspectives enhance inference accuracy, suggesting that additional viewpoints help capture richer causal structures in such settings.
Comparing virtual and realistic backgrounds, we find that realistic backgrounds introduce additional noise, making causal discovery more challenging. Our experiments indicate that even when models leverage prior real-world knowledge encoded in LLMs, the presence of realistic background information increases task complexity and negatively impacts inference performance (quantitative results shown in Fig. 39 in Appendix B).
Figure 8: We select two scenes, Spring and Parabola. Using the F1 score as metric, we assess VLM performance in causal discovery. The best and worst views are highlighted to demonstrate the impact of different perspectives. To analyze the effect of multi-view vs. single-view images, we average the performance across 9 individual views in each scene and compare it with the overall multi-view performance. More detailed version is in Appendix B.
4.5 Causal Intervention in VLMs
Evaluated Methods. To further assess the ability to learn causality from images, causal intervention serves as a crucial step beyond causal discovery. In this experiment, we leverage VLMs in interventions. This task is formulated as a multiple-choice problem, where the model must identify the correct outcomes based on induced causal changes. The detailed inference pipeline is illustrated in Fig. 9. In this setup, VLMs are provided with a few images along with questions about the variables present in the scene. By posing queries such as “What will happen after changing variable X?”, we expect the model to return the correct answer based on the causal relations depicted in the given images. This evaluation determines whether VLMs genuinely comprehend and reason causal relations rather than relying solely on statistical patterns.
In this experiment, intervention performance is evaluated with the accuracy of model selections. We assess 3 popular closed-source VLMs for this intervention task: ChatGPT-4o, Gemini-1.5-Pro, and Claude-3.5-Haiku.
Figure 9: Example of intervention prompts for trained VLMs.
Results and Analysis. As shown in Tab. 2, our experiments indicate that current VLMs struggle to handle complex rules, such as magnetic fields, in causal intervention tasks involving 3D scene images. We found that in this setting, the models’ inference results are predominantly influenced by the prior knowledge embedded within the LLMs, while visual cues—despite containing the ground truth—are largely ignored (shown in Fig. 10).
Figure 10: Case studies of failures. VLMs fail to grasp causality within physically consistent scenes for intervention tasks. In the failure case shown, VLMs overlook direct visual cues (indicated by the green frame), highlighting their limitations in causal reasoning based on visual information, showing a fundamental weakness in current VLMs when interpreting causal relations from images.
@@ 5 Conclusion
We introduce Causal3D, a comprehensive benchmark designed to evaluate causal reasoning in visual AI, including causal tasks of discovery, disentanglement, and intervention. Our dataset integrates structured causal graphs with corresponding 3D visual representations, providing a rigorous assessment framework across diverse physical and hypothetical scenes. Experimental results demonstrate that: First, the performance of current causal discovery algorithms decreases as the increasing of the complexity of causal structures. Second, for causal representation learning, the current state-of-arts can not well handle our realistic and diverse 3D scenes. Third, commercial VLMs struggle with causal inference based on visual cues and complex scenarios, particularly in hypothetical settings where prior knowledge is absent.
Causal3D serves as a critical step toward bridging the gap between traditional causal research and computer vision, enabling a more comprehensive and fine-grained evaluation of causal inference. We envision Causal3D as a foundation for future research, fostering advancements in causal-aware AI models and driving progress toward more reliable and interpretable machine intelligence.
Future Work. In the future, we plan to expand Causal3D dataset to encompass more specific and complex scenarios. Additionally, leveraging this dataset, we can fine-tune or further pre-train VLMs, enabling them to accurately learn and recognize causal relations within images. This approach will also enhance VLMs’ reasoning capabilities and improve their explainability.
Broader Impact
@@
Appendix A Additional Scenes and Data Details
Fig. 11, Fig. 12, and Fig. 13 show the dataset details of different scenes, including the following information: name of scenes, the number of causal variables (i.e., nodes in causal graph), the relation types (linear or non-linear), causal graphs, and structural equations. For realistic scenes, we also add a brief description. Furthermore, we showcase all the scenes by randomly sampling 2D images from different viewpoints and surroundings for illustration (shown from Fig. 14 to Fig. 38).

### v2

@@ 4.4 Causal discovery from few images
Evaluated Methods. Unlike traditional causal discovery methods that rely on tabular data or large-scale training images, we leverage pretrained VLMs to perform causal discovery using a small number of images. In this section we use ChatGPT [^2] https://platform.openai.com/docs/api-reference/introduction, Gemini [^3] https://ai.google.dev/api?lang=python and Claude [^4] https://www.anthropic.com/api to discover causal relations by few image examples and textual prompts in real and hypothetical scenes. The VLMs are tasked with uncovering causal relations by generating adjacency matrices representing causal graphs. Each prompt specifies key variables in the scenes, explicitly guiding the models to infer causal structures. We use 4 different prompting strategies, as detailed in Tab. 3 in the Appendix 7.
(a) Real-world scenes
(b) Hypothetical scenes.
Figure 5: Causal discovery results of VLMs in various scenes, averaged over datasets with 2–5 nodes in the causal graph.
1) Basic Prompts: General instructions that broadly guide the models to identify causal relations.
2) Explicit Function Prompts: The model is designated as a causal discovery expert to identify causal relationships among image variables.
@@
4) Few-Shot: The model is given three exemplar cases of causal discovery before performing the task, enabling it to generalize better by leveraging prior examples to improve causal relation identification. Performance is evaluated by comparing the generated adjacency matrices with the ground truth, providing a quantitative measure of the causal discovery task.
Results and Analysis. Demonstrated in Fig. 5(a), models performed significantly better in real-world scenes (e.g., governed by Hooke’s Law, magnetic fields and etc.) compared to hypothetical scenes, benefiting from prior physical knowledge embedded in their LLM components. However, as causal relations became more complex and involved more variables, performance declined, highlighting the challenge of uncovering causality in such scenes. In the hypothetical scenes, the results, as shown in Fig. 5(b), reveal poor performance. This suggests that when prior real-world knowledge is unavailable and only a limited number of images are provided, models consistently fail to uncover latent causal relations. Consequently, current closed-source VLMs have been shown to be unreliable for causal discovery in hypothetical settings.
Figure 6: We select two scenes, Spring and Parabola. Using the F1 score as metric, we assess VLM performance in causal discovery. The best and worst views are highlighted to demonstrate the impact of different perspectives. To analyze the effect of multi-view vs. single-view images, we average the performance across 9 individual views in each scene and compare it with the overall multi-view performance. More detailed version is in Appendix 7.
Views and Backgrounds. Causal3D provides multi-view images of the same scene, simulating real-world 3D environments with both virtual and realistic backgrounds. This enables in-depth analysis of how different viewpoints and background contexts affect causal discovery performance. Case studies from the Spring and Parabola scenes are shown in Fig. 6. Within the experiments, different 3D views affect the performance in causal discovery. Interestingly, our results reveal that intuitive views, such as a front view, are not always the most effective for uncovering latent causality. Different viewpoints can either increase or decrease task difficulty, with no single “golden standard” view for causal discovery. Additionally, when comparing multi-view and single-view inputs in the causal discovery task, multi-view performance varies depending on the scene (as shown in Fig. 9). In scenes with simple causal relations (e.g., a spring system with a linear relation among three variables), multi-view inputs tend to degrade model performance, possibly introducing unnecessary noise. Conversely, in more complex 3D scenes with intricate causal dependencies, like a parabola scene involving nonlinear relations among four variables in the virtual background, multi-view perspectives enhance inference accuracy, suggesting that additional viewpoints help capture richer causal structures in such settings.
Comparing virtual and realistic backgrounds, we find that realistic backgrounds introduce additional noise, making causal discovery more challenging. Our experiments indicate that even when models leverage prior real-world knowledge encoded in LLMs, the presence of realistic background information increases task complexity and negatively impacts inference performance (quantitative results shown in Fig. 37 in Appendix 7).
4.5 Causal intervention in VLMs
Evaluated Methods. To further assess the ability to learn causality from images, causal intervention serves as a crucial step beyond causal discovery. In this experiment, we leverage VLMs in interventions. This task is formulated as a multiple-choice problem, where the model must identify the correct outcomes based on induced causal changes. The detailed inference pipeline is illustrated in Fig. 8. In this setup, VLMs are provided with a few images along with questions about the variables present in the scene. By posing queries such as “What will happen after changing variable X?”, we expect the model to return the correct answer based on the causal relations depicted in the given images. This evaluation determines whether VLMs genuinely comprehend and reason causal relations rather than relying solely on statistical patterns. In this experiment, intervention performance is measured by model selection accuracy. We evaluate 3 popular commercial VLMs on this task: ChatGPT-4o, Gemini-1.5-Pro, and Claude-3.5-Haiku [^5] Model versions are those publicly available as of January 2025..
Figure 7: Example of intervention prompts for trained VLMs.
Figure 8: Case studies of failures. VLMs fail to grasp causality within physically consistent scenes for intervention tasks.
Table 2: Causal intervention in VLMs: Evaluation using three real-world rules, with accuracy (%) reported.
Models Reflection Lens Mag. Field
@@ 4.5 Causal intervention in VLMs
Results and Analysis. As shown in Tab. 2, our experiments indicate that current VLMs struggle to handle complex rules, e.g. magnetic fields, in causal intervention tasks. We found that the models’ inferences are dominated by the prior knowledge embedded in the LLMs, while visual cues—despite containing the ground truth—are largely ignored (see Fig. 8).
@@ 5 Conclusion
Causal3D is a comprehensive benchmark designed to evaluate causal reasoning in visual AI, including causal tasks of discovery, disentanglement, and intervention. Our dataset integrates structured causal graphs with corresponding 3D visual representations, providing a rigorous assessment framework across diverse physical and hypothetical scenes. Experimental results demonstrate that: 1) the performance of current causal discovery algorithms decreases as the increasing of the complexity of causal structures; 2) for causal representation learning, the current state-of-arts can not well handle our realistic and diverse 3D scenes; 3) commercial VLMs struggle with causal inference based on visual cues and complex scenarios, particularly in hypothetical settings where prior knowledge is absent. Causal3D serves as a critical step toward bridging the gap between traditional causal research and computer vision, enabling a more comprehensive and fine-grained evaluation of causal inference. We envision Causal3D as a foundation for future research, fostering advancements in causal-aware AI models and driving progress toward more reliable and interpretable machine intelligence.
Ethics Statement
This research adheres to the ICLR Code of Ethics. The study does not involve human subjects, personally identifiable information, or sensitive data. All datasets used are publicly available, and appropriate steps were taken to ensure compliance with privacy, fairness, and research integrity standards. No conflicts of interest or ethical concerns are anticipated.
Reproducibility Statement
We conducted extensive experiments with multiple random seeds to ensure the robustness and reproducibility of our results. Detailed descriptions of model architectures, training procedures, and hyperparameters closely follow those reported in the original baseline papers. The data generation code is submitted in a zip file as supplementary materials.
@@
6 Additional Scenes and Data Details
Fig. 9, Fig. 10, and Fig. 11 show the dataset details of different scenes, including the following information: name of scenes, the number of causal variables (i.e., nodes in causal graph), the relation types (linear or non-linear), causal graphs, and structural equations. For realistic scenes, we also add a brief description. Furthermore, we showcase all the scenes by randomly sampling 2D images from different viewpoints and surroundings for illustration (shown from Fig. 12 to Fig. 36). For notational simplicity, we omit the exogenous noise variables in the structural equations, all of which follow uniform distributions. For example, in the Reflection scene, the expression $A=B$ is a shorthand for the structural equation $B=A+\epsilon$, where $\epsilon$ is an independent uniform noise term. This non-Gaussianity ensures the identifiability of the causal direction $A\rightarrow B$.
@@ 7 Experiment Details
8 Broader Impact
9 Limitations
Although Causal3D is comprehensive, there remains room for improvement. First, the current benchmark is constructed solely from observational data; incorporating interventional data in the future would enable user interaction and support more interactive causal evaluation. Second, the dataset complexity could be further enriched by introducing more nodes in the causal graphs and incorporating finer-grained visual details, such as textures of objects.
10 The Use of Large Language Models (LLMs)
We used ChatGPT, Claude and Gemini to evaluate our datasets. The LLM was not involved in data generation, model training, or writing of technical content.